Consider the parametric curve described by r(t) =〈2 sin(3t), √4t, 2cos(3t)〉. Find the unit tangent vector of r(t).
1
Expert's answer
2020-05-27T16:46:07-0400

The unit tangent vector to the parametric curve "r(t)" is given by,


"\overrightarrow{T}(t)=\frac{\overrightarrow{r'}(t)}{∥ \overrightarrow{r'}(t)∥}"


Here,


"\overrightarrow{r'}(t)=⟨ 2\frac{d}{dt}(\sin(3t)),\frac{d}{dt}(\sqrt{4}t),2\frac{d}{dt}(\cos(3t))⟩"


"=⟨ 2(3\cos(3t)),\sqrt{4},2(-3\sin(3t))⟩"


"=⟨ 6\cos(3t),2,-6\sin(3t)⟩"


"∥\overrightarrow{r'}(t)∥=\sqrt{(6\cos(3t))^2+(\sqrt{4})^2+(-6\sin(3t))^2}"


"=\sqrt{36\cos^2(3t)+4+36\sin^2(3t)}"


"=\sqrt{36(\cos^2(3t)+\sin^2(3t))+4}"


Use trigonometric identity "\cos^2(x)+sin^2(x)=1" to simplify the term inside the radical as,


"=\sqrt{36(1)+4}"


"=\sqrt{36+4}"


"=\sqrt{40}"


"=2\sqrt{10}"


Therefore, the unit tangent vector is,


"\overrightarrow{T}(t)=\frac{⟨ 6\cos(3t),2,-6\sin(3t)⟩}{∥ ⟨ 6\cos(3t),\sqrt{4},-6\sin(3t)⟩∥}"


"=\frac{⟨ 6\cos(3t),2,-6\sin(3t)⟩}{2\sqrt{10}}"


"=⟨ \frac{3\cos(3t)}{\sqrt{10}},\frac{1}{\sqrt{10}},-\frac{3\sin(3t)}{\sqrt{10}}⟩"


Therefore, the unit tangent vector to the parametric curve is "\overrightarrow{T}(t)=⟨ \frac{3\cos(3t)}{\sqrt{10}},\frac{1}{\sqrt{10}},-\frac{3\sin(3t)}{\sqrt{10}}⟩".